OpenStudy (anonymous):

(2, -1) is the solution to this system of equations. -2y=x 2y=x-4 true or false?

OpenStudy (anonymous):

true

OpenStudy (anonymous):

PLUG (2, -1) BACK INTO THE SYSTEM OF EQUATIONS. IF IT IS TRUE FOR BOTH THEN YES IT IS A SOLTUTION.

OpenStudy (anonymous):

(2, -1) -2y=x 2y=x-4 -2(-1) = 2 2(-1) = 2-4 2 = 2 -2 = 2-4 2 = 2 -----> TRUE -2 = -2 -------> TRUE